Ticket Number: HEA-NOC/20080202-3 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver:
Ticket Opened: 20080202 22:37 UTC Problem Start: 20080202 20:02 UTC
Ticket Update: 20080203 13:39 UTC Problem End:
Loss of connectivity to Hermes Schools
No access to/from Hermes Schools
20080202 22:27 DL Systems report site inaccessible since 20:02 today.
Contact Service Provider to communicate circumstance.
20080203 13:00 DL Call Service Provider and leave message in voicemail
asking for update.
20080203 13:36 DL Receive call from Service Provider notifying the problem
is recurrent, usually during weekends and related to other equipment in the
are using the same frequency (this a wireless link) than the transmission
system involved in this ticket. Equipment has been rebooted at 13:00 as
part of troubleshooting but that has not solved the problem. The Service
Provider has not other means of troubleshooting at this point in time and
has to escalate to third level support during working weekdays to continue
Time to Fix:
Unknown at the moment
This ticket can be monitored at http://www.hea.net/tickets/20080202-3
HEAnet Limited. Registered in Ireland, No. 275301.